calculate maximum amount of product produced and amount of any
unreacted excess reagent, knowing the limiting reagent in a reaction
1
Expert's answer
2016-09-02T09:59:03-0400
Assuming that in the process of carbon burning, carbon is being consumed fully:
C+O2=CO2
The amount of carbon available – 6 g (limiting reactant).

The amount of oxygen, required for its burning is:
n (C)=n(O2)
n=m/M
M(C)=12 g/mol
M(O2)=32 g/mol
n(C)=6/12=0.5 mol
m (O2)=0.5·32=16 g

The amount of product (CO2) formed is:
n (C)=n(CO2)
n=m/M
M(C)=12 g/mol
M(CO2)=44 g/mol
m (CO2)=0.5·44=22 g